A 5ft diameter solid wheel initially rotating at 50.0rad/s comes to a uniform stop after 8.0 rotations in radians. What is the tangential acceleration at the rim of the wheel?


1
Expert's answer
2021-12-12T16:40:01-0500

The tangential acceleration is


"a_\tau=\alpha r=\frac{\omega^2}{2\phi} ·\frac d2=390.6\text{ ft/s}^2."